Write each parabola in standard form and state vertex line of symmetry and which way it opens
1. 5x(squared)=Y
2. X-1=(Y+2)(squared)+3
3. X=3Y(Squared)+14

Answer by lwsshak3(11628)

Write each parabola in standard form and state vertex line of symmetry and which way it opens
1. 5x(squared)=Y
2. X-1=(Y+2)(squared)+3
3. X=3Y(Squared)+14
..
1. 5x(squared)=Y
Standard form: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex. Parabola opens upwards if A is positive and downward if A is negative. A is just a multiplier that affects the steepness or narrowness of the parabola. Larger A makes the curve steeper.
For given equation: y=5(x-0)^2+0 (rewrite)
Vertex (0,0)
Line of symmetry x=0
Opens upwards
..
2. X-1=(Y+2)(squared)+3
Standard form:x=A(y-k)^2+h, with (h,k) being the (x,y) coordinates of the vertex. Parabola opens rightwards if A is positive and leftward if A is negative. A is just a multiplier that affects the width or narrowness of the parabola. Larger A makes the curve narrower.
For given equation:x=(y+2)^2+4 (rewrite)
vertex (4,2)
Line of symmetry y=2
Opens rightwards
..
3. X=3Y(Squared)+14
Standard Form: same as that of problem 2 above.
For given equation: x=3(y-0)^2+14 (rewrite)
Vertex (14,0)
Line of symmetry y=0 or x-axis
Opens rightwards